# Symmetries of the periodic Fredkin chain

### v1

@@ Abstract.
Fredkin chain is a spin-$1/2$ model with interaction of three nearest neighbors. In the case of periodic boundary conditions the ground state is degenerated and can be described in terms of equivalence classes of the Dyck paths. We introduce two operators commuting with the Hamiltonian, which play the roles of lowering and raising operators when acting at the ground states. These operators generate the $B$- or $C$-type Lie algebras, depending on whether the number of sites $N$ is odd or even, respectively, with rank $n=\lceil N/2\rceil$. The third component of total spin operator can be represented as a sum of the Cartan subalgebra elements and some central element. In the $C$-type Lie algebra case (even number of sites) this representation coincides with a similar formula previously conjectured for spin-$1$ operators, in the context of periodic Motzkin chain.
@@ 1. Introduction
The Fredkin gate was originally introduced in the context of reversible quantum calculations [FT-82]. It is also known as controlled-SWAP gate and involves three qubits in operation. Relatively recently it was experimentally realized with the help of entangled photons [PHFRP-16].
@@
The Fredkin chain is a model of interacting $s=1/2$ spins generalizing in a certain sense the famous Heisenberg XXX chain [SK-17, SK-18]. In this model, the spin-00 component of two $s=1/2$ spins interacts with the third neighbor (left or right). Under suitably chosen open boundary conditions the ground state of the Hamiltonian can be given as a homogeneous sum of Dyck paths. This state is highly entangled: it shows a logarithmic growth of entanglement entropy with respect to the size of a subsystem. The similar phenomenon is known in the context Motzkin chain, which is model of nearest-neighbor interacting $s=1$ spins, and with the ground state given by a sum of Motzkin paths [BCMNS-12, M-17].
@@
Furthermore, the Fredkin chain appears to be even more related to the Motzkin chain, as it can be viewed as an half-integer spin analogue of the latter. Both models exhibit criticality without a “frustration”: the energy gap vanishes for large system sizes and the unique ground states minimize all individual terms of the Hamiltonians. This property ensures stability of the ground state against inclusion in the Hamiltonian term-dependent interaction parameters. The similarities extend to “colored” versions of these models corresponding to their higher spin generalizations [MS-16, SUZKK-17, ZAK-17]. For these reasons both models are of crucial importance for study of long-range entanglement in critical one-dimensional quantum systems and related topics [DaBT-19, AAZK-19, SVG-23, CBG-24, MGM-24, WZYK-24, VPKGRP-24, CSGV-25].
@@
An interesting question about the Fredkin chain to be raised whether it is an quantum integrable model. The Heisenberg XXX chain is a famous paradigmatic example of such a system [TF-79, S-82, TTF-83, KBI-93, F-96, S-20]. Integrability has been found for a “free” version of the Motzkin chain, with one of the terms describing interaction of spins removed [HSK-23]. Recently, it was observed, albeit purely experimentally for systems of small sizes, that the Motzkin chain with periodic boundary conditions possesses a quite rich symmetry algebra [P-25], that could be a signal of its quantum integrability.
@@
The purpose of the present study is to obtain similar results for the periodic Fredkin chain. Alike for the Motzkin chain, our study here is mainly conjectural, although the central result about properties of the raising and lowering operators is rigorous (see Theorem 2 below). Other results are collected in four conjectures formulated by studying systems of small size, for systems of up to eight sites in the length.
@@
Our main object which we study are operators which play the roles of lowering and raising operators when acting the ground states. These operators decreasing and increasing, respectively, the eigenvalue of the third component of total spin operator, which labels these states. In the route of our study we have also encountered an alternative though less explicit representation for these operators in which they are expressed in terms other components of the total spin operator. This immediately implies that the same operators also commute with the Hamiltonian of the Heisenberg XXX spin-$1/2$ chain. On the other hand, we also conjecture that these operators generate the $B$- or $C$-type Lie algebras, depending on whether the number of sites $N$ is odd or even, respectively, with rank $n=\lceil N/2\rceil$. Furthermore, there exists a central element of the symmetry algebra which can be given as a sum the third component of total spin operator and Cartan subalgebra elements. Surprisingly, we find that in the $C$-type case (even number of sites) this representation coincides with a similar formula previously appeared in the context of periodic Motzkin chain, that hints at the fact that this relation is somewhat representation independent.
@@
The paper is organized as follows. In the next section we review known basic results about the periodic Fredkin chain. In section 3 we introduce the raising and lowering operators and the main result about their properties. In section 4 we study the algebra they generate and show that the third component of total spin operator contains a central element extending this algebra. We conclude in section 5 with a brief discussion of the results and possible proofs of the formulated conjectures.
@@ 2. The Hamiltonian, Dyck paths, and the ground states
In this section we recall main ingredients from [SK-17] about the Fredkin chain, with a focus on the periodic boundary conditions.
@@ 2.1. The Hamiltonian
We will use also projectors to the spin-up and spin-down states:
@@
The space of states of the $N$-site Fredkin chain is the vector space $(\mathbb{C}^{2})^{\otimes N}$ in which the $j$th factor corresponds to the $j$th site of the spin chain. The basis vectors of $(\mathbb{C}^{2})^{\otimes N}$ are denoted by
@@
and by $\sigma_{j}^{\pm,z}$ — operators acting as Pauli matrices in the $j$th copy of $(\mathbb{C}^{2})^{\otimes N}$. For a later reference we need the “total spin operator” for which we choose different normalization of components,
@@
such that the eigenvalues of $\mathrm{S}^{z}$ are $S^{z}=-\frac{N}{2},-\frac{N}{2}+1,\dots,\frac{N}{2}-1,\frac{N}{2}$ (we follow the convention to use upright capital letters for “global” operators and their italic variants for their eigenvalues).
Consider two adjacent copies of $\mathbb{C}^{2}$, say, $j$th and $(j+1)$th. Important objects are the permutation operator $\mathrm{P}_{j,j+1}$ which exchanges these copies, and the projector $\Pi_{j,j+1}$ to the spin-00 state, which is essentially the antisymmetrizer,
@@
The Hamiltonian density of the Fredkin spin chain describes an interaction the spin-00 component of two $s=1/2$ spins with a neighboring third such spin as follows:
@@
The periodic boundary conditions mean that the Hamiltonian is the cyclic sum
@@
$\mathrm{C}=\mathrm{P}_{1,2}\cdots\mathrm{P}_{N-2,N-1}\mathrm{P}_{N-1,N}.$ (2.9)
is the cyclic shift operator $\mathrm{C}$ acting at sites as $i\mapsto i+1$, and it is an integral of motion,
@@
The operator $\mathrm{S}^{z}$ is also an integral of motion
$\left[\mathrm{S}^{z},\mathrm{H}\right]=0,$ (2.11)
thus eigenvalues of both the cyclic shift operator $\mathrm{C}$ and the third component of total spin operator $\mathrm{S}^{z}$ can be used to label eigenstates of the Hamiltonian.
@@ 2.2. Dyck paths
We only summarize the construction developed in [SK-17] to the extent which is sufficient to describe ground states of the Hamiltonian (2.8) below.
@@
Apparently, all possible Dyck paths of length $N$ are in a one-to-one correspondence with the basis vectors of $(\mathbb{C}^{2})^{\otimes N}$, the state space of the Fredkin chain. The correspondence is based on a simple observation that for each letter $\ell_{j}$ in (2.4) one can associate step up or step down according to whether $\ell_{j}=\uparrow$ or $\ell_{j}=\downarrow$. Thus, a string (or, a “word”) $\ell_{1}\dots\ell_{N}$ uniquely defines the Dyck path, and this can be used for their notation.
Division of Dyck paths on equivalence classes introduces certain order among the basis vectors. In particular, vectors associated with the class $C_{a,b}(N)$ are eigenvectors of $\mathrm{S}^{z}$ with the eigenvalue $S^{z}=\frac{1}{2}(b-a)$. Furthermore, one can introduce states $\left\lvert C_{a,b}(N)\right\rangle$ as homogeneous sums over Dyck paths belonging to $C_{a,b}(N)$, namely,
@@
and there are also the states $\left\lvert C_{3,0}(3)\right\rangle$, $\left\lvert C_{1,0}(3)\right\rangle$, and $\left\lvert C_{2,1}(3)\right\rangle$ obtained by reversal of arrows of these states (mirror transformation of paths), respectively.
@@
The Dyck paths for $N$ even belonging to the class $C_{0,0}(N)$ correspond to the Dyck paths in their standard definition, i.e., these are paths that start at $(0,0)$ and ends at $(0,N)$ with the condition that they are allowed to pass below the $x$-axis. The corresponding state $\left\lvert C_{0,0}(N)\right\rangle$ is a unique ground state of the open Fredkin chain introduced in [SK-17, SK-18].
@@ 2.3. Ground states
In the case of periodic boundary conditions, the ground states corresponding to zero eigenvalue of the Hamiltonian can be constructed as superposition of states $\left\lvert C_{a,b}(N)\right\rangle$ with $a$ and $b$ related to an eigenvalue of the operator $\mathrm{S}^{z}$. A key observation leading to expressions for the ground states is an extension of the Fredkin moves by periodicity, see [SK-17, SK-18], for details. To emphasize that this an established result, we formulate it here as a theorem.
@@
For each value of $S^{z}\in\{-\frac{N}{2},-\frac{N}{2}+1,\dots,\frac{N}{2}-1,-\frac{N}{2}\}$ there exists one and only one ground state $\left\lvert v_{S^{z}}\right\rangle$ of the periodic Fredkin chain Hamiltonian (2.8) with zero eigenvalue, which correspond to the eigenvalue $C=1$ of the cyclic shift operator, of the form
@@
In the case of $N$ even, $N=2n$, and $S^{z}=0$ there exists one more ground state which correspond to the eigenvalue $C=-1$, of the form
@@
Thus [^1] In [SK-17, SK-18] the degeneracies for odd and even $N$ were misprinted as $N$ and $N+1$, respectively., in the case of $N$ odd the degeneracy of the ground state is $N+1$ and in the case of $N$ even is $N+2$.
@@
Appearance of two ground states for $N$ even and $S^{z}=0$ comes from the fact that the considerations involving the periodic Fredkin moves lead to existence of two eigenstates, with even and odd values of $a$, for the classes $C_{a,a}(N)$. An example of $N=4$ is shown in Figure 3. These states are
@@ 3. Lowering and raising operators
Now we turn to the results. We have obtained them mainly by a direct study of the model for up to $N=8$. The central result, about explicit expressions for raising and lowering operators, is given here as a theorem. The remaining results (in this and the next section) are summarized in four conjectures.
@@ 3.1. Symmetry generators
In [SK-17, SK-18] it was indicated that the states $\left\lvert v_{S^{z}}\right\rangle$ are nothing but the ground states of the spin-$1/2$ ferromagnetic Heisenberg XXX chain (see, e.g., [F-96]). As such, they can be obtained from each other by acting with the operators $\mathrm{S}^{+}$ and $\mathrm{S}^{-}$, which increase and decrease, respectively, the value of $S^{z}$. However, these operators has no similar role in the context of the Fredkin chain since they do not commute with the Hamiltonian (2.8).
One of our main results here is that there exists another pair of operators which play the similar role of raising and lowering operators when acting at the states $\left\lvert v_{S^{z}}\right\rangle$, but commuting with the Hamiltonian (2.8).
@@
when acting at the ground states $\left\lvert v_{S^{z}}\right\rangle$ play the role of raising and lowering operators, respectively,
@@
where $c_{\pm}(S^{z})\neq 0$ are some constants, and commute with the Hamiltonian,
@@ 3.2. A relation to the total spin operator
As it clear from (3.1), the matrices representing operators $\Sigma^{\pm}$ are build of 00’s and $1$’s; these matrices are nilpotent, $(\Sigma^{\pm})^{N+1}=0$, though not strictly upper or lower triangle ones. The operators $\Sigma^{\pm}$ are non-local, contrary to the operators $\mathrm{S}^{\pm}$, and at a first glance, they are unrelated. However, it turn out that $\Sigma^{\pm}$’s can be expressed in terms of $\mathrm{S}^{\pm}$’s.
@@
where $\gamma_{k}$ are some coefficients and $\operatorname*{\mathrm{aad}}$ denotes the anti-adjoint action, $\operatorname*{\mathrm{aad}}a\equiv\{a,\cdot\,{}\}$, where braces stand for the anti-commutator.
For example, for $N=3$,
@@
Coefficients $\gamma_{k}$ for the cases of $N$ up to $N=8$ are
$\begin{gathered}\gamma_{1}=-\frac{1}{4},\quad\gamma_{2}=\frac{1}{8},\qquad N=3,\\ \gamma_{1}=-\frac{3}{4},\quad\gamma_{2}=\frac{1}{8},\qquad N=4,\\ \gamma_{1}=\frac{1}{8},\quad\gamma_{2}=-\frac{7}{96},\quad\gamma_{3}=\frac{1}{192},\qquad N=5,\\ \gamma_{1}=\frac{5}{8},\quad\gamma_{2}=-\frac{13}{96},\quad\gamma_{3}=\frac{1}{192},\qquad N=6,\\ \gamma_{1}=-\frac{5}{64},\quad\gamma_{2}=\frac{37}{768},\quad\gamma_{3}=-\frac{11}{2304},\quad\gamma_{4}=\frac{1}{9216},\qquad N=7,\\ \gamma_{1}=-\frac{35}{64},\quad\gamma_{2}=\frac{101}{768},\quad\gamma_{3}=-\frac{17}{2304},\quad\gamma_{4}=\frac{1}{9216},\qquad N=8.\end{gathered}$ (3.8)
The values of $\gamma_{1}$ at $N=7$ and $N=8$ breaks an easy guess that they should differ be $1/2$, as it happens at $N=3$ and $N=4$, and at $N=5$ and $N=6$.

### v2

@@ Abstract.
The Fredkin chain is a spin-$1/2$ model with interaction of three nearest neighbors. In the case of periodic boundary conditions, the ground state is degenerate and can be described in terms of equivalence classes of Dyck paths. We introduce two operators commuting with the Hamiltonian which play the roles of lowering and raising operators when acting on the ground states. These operators generate the $B$- or $C$-type Lie algebras, depending on whether the number of sites $N$ is odd or even, respectively, with rank $n=\lceil N/2\rceil$. The third component of the total spin operator can be represented as a sum of the Cartan subalgebra elements and some central element. In the $C$-type Lie algebra case (even number of sites), this representation coincides with a similar formula previously conjectured for spin-$1$ operators, in the context of the periodic Motzkin chain.
@@ 1. Introduction
The Fredkin gate was originally introduced in the context of reversible quantum computation [FT-82]. It is also known as the controlled-SWAP gate and involves three qubits in its operation. Relatively recently, it was experimentally realized with the help of entangled photons [PHFRP-16].
@@
The Fredkin chain is a model of interacting $s=1/2$ spins, generalizing in a certain sense the famous Heisenberg XXX chain [SK-17, SK-18]. In this model, the spin-0 component of two $s=1/2$ spins interacts with the third neighbor (left or right). Under suitably chosen open boundary conditions, the ground state of the Hamiltonian can be given as a homogeneous sum of Dyck paths. This state is highly entangled: it shows a logarithmic growth of entanglement entropy with respect to the size of a subsystem. A similar phenomenon is known in the context of the Motzkin chain, which is model of nearest-neighbor interacting $s=1$ spins, and with the ground state given by a sum of Motzkin paths [BCMNS-12, M-17].
@@
Furthermore, the Fredkin chain appears to be even more related to the Motzkin chain, as it can be viewed as a half-integer spin analogue of the latter. Both models exhibit criticality without “frustration”: the energy gap vanishes for large system sizes, and the unique ground states minimize all individual terms of the Hamiltonians. This property ensures stability of the ground state against inclusion in the Hamiltonian of term-dependent interaction parameters. The similarities extend to arbitrary half-integer and integer spin versions of these models, referred to as “colored” Fredkin and Motzkin chains, respectively [MS-16, SUZKK-17, ZAK-17]. Further generalizations include the construction of various systems such as tensor networks [AAZK-19] and a solvable cellular automaton [SVG-23]. Specifically, the Fredkin chain and the Dyck paths description of its ground state can be connected to complex multi-matrix models [WZYK-24] and Moore–Read quantum Hall states [VPKGRP-24]. For results on the study of correlations, dynamics, and entanglement in Fredkin and Motzkin spin chains, see, e.g., [DaBT-19, CBG-24, MGM-24] and references therein. Recently, it was shown [CSGV-25] that both models possess a pseudo-local conserved charge, and transport in these systems is subdiffusive.
@@
An interesting question to raise about the Fredkin chain is whether it is a quantum integrable model. Integrability of a spin chain implies that there exist Yang–Baxter structures that guarantee the existence of a hierarchy of conserved quantum integrals of motion [TF-79, S-82, TTF-83]. A paradigmatic example of such a system is the Heisenberg XXX chain; the Fredkin chain can be seen as its generalization. For a review of the methods exploiting integrability, see, e.g., the monograph [KBI-93] and lecture notes [F-96]; a contemporary exposition can be found in [S-20]. Integrability has been found for the “free” version of the Motzkin chain, in which one of the terms describing the interaction of spins is removed from the Hamiltonian [HSK-23]. Recently, it was observed, albeit for systems of small sizes, that the Motzkin chain with periodic boundary conditions possesses a quite rich symmetry algebra [P-25], that could be a signal of its quantum integrability.
@@
The purpose of the present study is to obtain similar results for the periodic Fredkin chain. As for the Motzkin chain, our study here is mainly conjectural, although the central result about the properties of the raising and lowering operators is rigorous (see Theorem 2 below). Other results are collected in four conjectures, formulated by studying systems of small size, namely, the chains with $N\leqslant 10$ sites.
@@
The main object we study is the set of two operators which play the roles of lowering and raising operators when acting on the ground states. These operators decrease and increase, respectively, the eigenvalue of the third component of the total spin operator, which labels these states. In the route of our study, we have also encountered an alternative, though less explicit, representation for these operators, in which they are expressed in terms of other components of the total spin operator. This immediately implies that the same operators also commute with the Hamiltonian of the Heisenberg XXX spin-$1/2$ chain. We conjecture that these operators generate the $B$- or $C$-type Lie algebras, depending on whether the number of sites $N$ is odd or even, respectively, with rank $n=\lceil N/2\rceil$. Furthermore, there exists a central element of the symmetry algebra, which can be given as a sum of the third component of the total spin operator and Cartan subalgebra elements. Surprisingly, we find that in the $C$-type case (even number of sites) this representation coincides with a similar formula that previously appeared in the context of the periodic Motzkin chain. This observation suggests that the connection between the third component of the total spin operator, Cartan subalgebra elements, and the central element is independent of the $\mathfrak{sl}_{2}$ algebra representation.
@@
The paper is organized as follows. In the next section, we review known basic results about the periodic Fredkin chain. In Section 3 we introduce the raising and lowering operators and prove the main result about their properties. In Section 4 we study the algebra they generate and show that the third component of the total spin operator contains a central element extending this algebra. We conclude in Section 5 with a brief discussion of the results and possible proofs of the formulated conjectures.
@@ 2. The Hamiltonian, Dyck paths, and the ground states
In this section, we recall the main ingredients from [SK-17] about the Fredkin chain, with a focus on periodic boundary conditions.
@@ 2.1. The Hamiltonian
We will also use projectors onto the spin-up and spin-down states:
@@
The space of states of the $N$-site Fredkin chain is the vector space $(\mathbb{C}^{2})^{\otimes N}$, in which the $j$th factor corresponds to the $j$th site of the spin chain. The basis vectors of $(\mathbb{C}^{2})^{\otimes N}$ are denoted by
@@
and by $\sigma_{j}^{\pm,z}$ — operators acting as Pauli matrices in the $j$th copy of $(\mathbb{C}^{2})^{\otimes N}$. For later reference, we need the “total spin operator”, for which we choose a different normalization of its components,
@@
such that the eigenvalues of $\mathrm{S}^{z}$ are $S^{z}=-\frac{N}{2},-\frac{N}{2}+1,\dots,\frac{N}{2}-1,\frac{N}{2}$ (we follow the convention of using upright capital letters for “global” operators and their italic variants for their eigenvalues).
Consider two adjacent copies of $\mathbb{C}^{2}$, say, the $j$th and $(j+1)$th. Important objects are the permutation operator $\mathrm{P}_{j,j+1}$, which exchanges these copies, and the projector $\Pi_{j,j+1}$ onto the spin-0 state, which is essentially the antisymmetrizer,
@@
The Hamiltonian density of the Fredkin spin chain describes an interaction of the spin-0 component of two $s=1/2$ spins with a neighboring third such spin, as follows:
@@
The periodic boundary conditions mean that the Hamiltonian is the cyclic sum,
@@
$\mathrm{C}=\mathrm{P}_{1,2}\cdots\mathrm{P}_{N-2,N-1}\mathrm{P}_{N-1,N}$ (2.9)
is the cyclic shift operator $\mathrm{C}$, acting on sites as $i\mapsto i+1$, and it is an integral of motion,
@@
The operator $\mathrm{S}^{z}$ is also an integral of motion,
$\left[\mathrm{S}^{z},\mathrm{H}\right]=0.$ (2.11)
Thus, eigenvalues of both the cyclic shift operator $\mathrm{C}$ and the third component of the total spin operator $\mathrm{S}^{z}$ can be used to label eigenstates of the Hamiltonian.
@@ 2.2. Dyck paths
We only summarize the construction developed in [SK-17] to the extent that is sufficient to describe ground states of the Hamiltonian (2.8) below.
@@
Apparently, all possible Dyck paths of length $N$ are in one-to-one correspondence with the basis vectors of $(\mathbb{C}^{2})^{\otimes N}$, the state space of the Fredkin chain. The correspondence is based on a simple observation that for each letter $\ell_{j}$ in (2.4) one can associate step up or step down according to whether $\ell_{j}=\uparrow$ or $\ell_{j}=\downarrow$. Thus, a string (or, a “word”) $\ell_{1}\dots\ell_{N}$ uniquely defines the Dyck path, and this can be used for its notation.
Division of Dyck paths into equivalence classes introduces certain order among the basis vectors. In particular, vectors associated with the class $C_{a,b}(N)$ are eigenvectors of $\mathrm{S}^{z}$ with the eigenvalue $S^{z}=\frac{1}{2}(b-a)$. Furthermore, one can introduce states $\left\lvert C_{a,b}(N)\right\rangle$ as homogeneous sums over Dyck paths belonging to $C_{a,b}(N)$, namely,
@@
and there are also the states $\left\lvert C_{3,0}(3)\right\rangle$, $\left\lvert C_{1,0}(3)\right\rangle$, and $\left\lvert C_{2,1}(3)\right\rangle$ obtained by reversal of spins of these states (mirror transformation of paths), respectively.
@@
The Dyck paths for $N$ even belonging to the class $C_{0,0}(N)$ correspond to the Dyck paths in their standard definition, i.e., these are paths that start at $(x,y)=(0,0)$ and end at $(x,y)=(0,N)$ with the condition that they are not allowed to pass below the $x$-axis. The corresponding state $\left\lvert C_{0,0}(N)\right\rangle$ is the unique ground state of the open Fredkin chain introduced in [SK-17, SK-18].
@@ 2.3. Ground states
In the case of periodic boundary conditions, the ground states corresponding to zero eigenvalues of the Hamiltonian can be constructed as a superposition of states $\left\lvert C_{a,b}(N)\right\rangle$ with $a$ and $b$ related to an eigenvalue of the operator $\mathrm{S}^{z}$. A key observation leading to expressions for the ground states is an extension of the Fredkin moves by periodicity, see [SK-17, SK-18], for details. To emphasize that this is an established result, we formulate it here as a theorem.
@@
For each value of $S^{z}\in\{-\frac{N}{2},-\frac{N}{2}+1,\dots,\frac{N}{2}-1,\frac{N}{2}\}$ there exists one and only one ground state $\left\lvert v_{S^{z}}\right\rangle$ of the periodic Fredkin chain Hamiltonian (2.8) with the zero eigenvalue, which corresponds to the eigenvalue $C=1$ of the cyclic shift operator, of the form
@@
In the case of $N$ even, $N=2n$, and $S^{z}=0$ there exists one more ground state which corresponds to the eigenvalue C=1C=-1, of the form
@@
Thus [^1] In [SK-17, SK-18] the degeneracies for odd and even $N$ were misprinted as $N$ and $N+1$, respectively., in the case of $N$ odd the degeneracy of the ground state is $N+1$ and in the case of $N$ even, it is $N+2$.
@@
Appearance of two ground states for $N$ even and $S^{z}=0$ comes from the fact that the considerations involving the periodic Fredkin moves lead to existence of two eigenstates, with even and odd values of $a$ for the classes $C_{a,a}(N)$. An example of $N=4$ is shown in Figure 3. These states are
@@ 3. Lowering and raising operators
Now we turn to the results. We have obtained them mainly by a direct study of the model with $N\leqslant 10$ sites. The central result, about explicit expressions for raising and lowering operators, is given here as a theorem. The remaining results (in this and the next section) are summarized in four conjectures.
@@ 3.1. Symmetry generators
In [SK-17, SK-18] it was indicated that the states $\left\lvert v_{S^{z}}\right\rangle$ are nothing but the ground states of the spin-$1/2$ ferromagnetic Heisenberg XXX chain (see, e.g., [F-96]). As such, they can be obtained from each other by acting with the operators $\mathrm{S}^{+}$ and $\mathrm{S}^{-}$, which increase and decrease, respectively, the value of $S^{z}$. However, these operators have no similar role in the context of the Fredkin chain since they do not commute with the Hamiltonian (2.8).
One of our main results here is that there exists another pair of operators which play a similar role of raising and lowering operators when acting on the states $\left\lvert v_{S^{z}}\right\rangle$, but commuting with the Hamiltonian (2.8).
@@
when acting on the ground states $\left\lvert v_{S^{z}}\right\rangle$ play the role of raising and lowering operators, respectively,
@@
where $c_{\pm}(S^{z})\neq 0$ are some constants. These operators commute with the Hamiltonian,
@@ 3.2. A relation to the total spin operator
As it is clear from (3.1), the matrices representing operators $\Sigma^{\pm}$ are built of 0’s and $1$’s; these matrices are nilpotent, $(\Sigma^{\pm})^{N+1}=0$, though not strictly upper or lower triangular ones. The operators $\Sigma^{\pm}$ are non-local, contrary to the operators $\mathrm{S}^{\pm}$, and at a first glance they are unrelated. However, it turns out that $\Sigma^{\pm}$’s can be expressed in terms of $\mathrm{S}^{\pm}$’s.
@@
where $\gamma_{k}$ are some coefficients and $\operatorname*{\mathrm{aad}}$ denotes the anti-adjoint action, $(\operatorname*{\mathrm{aad}}a)\,b\equiv\{a,b\}=ab+ba$.
For example, in the case $N=3$ we find
@@
Coefficients $\gamma_{k}$ for cases $N\leqslant 10$ are given in Table 1.
Table 1. Coefficients $\gamma_{k}$
NN $\gamma_{1}$ $\gamma_{2}$ $\gamma_{3}$ $\gamma_{4}$ $\gamma_{5}$
3 $\frac{1}{4}$ $\frac{1}{8}$
4 $\frac{3}{4}$ $\frac{1}{8}$
5 $\frac{1}{8}$ $\frac{7}{96}$ $\frac{1}{192}$
6 $\frac{5}{8}$ $\frac{13}{96}$ $\frac{1}{192}$
7 $\frac{5}{64}$ $\frac{37}{768}$ $\frac{11}{2304}$ $\frac{1}{9216}$
8 $\frac{35}{64}$ $\frac{101}{768}$ $\frac{17}{2304}$ $\frac{1}{9216}$
9 $\frac{7}{128}$ 53315 360-$\frac${533}{15\,360} $\frac{727}{184\,320}$ 536 864-$\frac${5}{36\,864} $\frac{1}{737\,280}$
10 $\frac{63}{128}$ $\frac{641}{5120}$ $\frac{509}{61\,440}$ 736 864-$\frac${7}{36\,864} $\frac{1}{737\,280}$